Nalu and Kamaile take-out a mortgage in the amount of $585,000 to purchase an apartment as their principal place of residence. They are able to obtain a 30-year mortgage at a fixed rate of 6%. Below, you will be asked for the amount of their monthly payment and for a aggregated amount of interest that they paid.

What is the value of that factor?

What is their monthly payment?

n Outstanding Principal INTEREST Amount Excess PMT

1

2

What is there deduction for interest for this mortgage annuity in the year that it was issued?

I need the calculation and equation.
Answer #1

Loan Amount = $585,000

Time Period = 30 years

Interest Rate = 6%

Calculating Monthly Payment,

PMT = [FV = 0 , T = 360, I = 0.06/12, PV = 585,000]

PMT = $3,507.37

Interest Paid = 3,507.37(360) - 585,000

Interest Paid = $677,653.20